OpenStudy (anonymous):

Graph the Linear equation by solving for y first. 3x+5y=20 ... i got y to one side.. ended with y = =3/5x+4 .. i need help from here..

OpenStudy (amistre64):

divide off the 20, and then the tops; to show you your x and y intercepts

OpenStudy (anonymous):

Graphing help? http://www.wolframalpha.com/input/?i=3x%2B5y%3D20

OpenStudy (amistre64):

\[\frac{3x}{20}+\frac{5y}{20}=\frac{20}{20}\] \[\frac{x}{20/3}+\frac{y}{20/5}=1\] \[\frac{x}{20/3}+\frac{y}{4}=1\]

OpenStudy (amistre64):

make a point at 20/3, x and 0,4; then connect the dots :)

OpenStudy (amistre64):

you dont need to "first solve for y" in order to graph the line .... that step is just to get you into the feel of how the slope intercept form operates

OpenStudy (anonymous):

That is a pretty neat trick... I did not know that. Why does that work?

OpenStudy (amistre64):

it works because that is how the intercepts can be defined. in order for the solution to be 1 when x or y is zero; you have to have the same top as the bottom right?

OpenStudy (amistre64):

hypers and ellipses are the same way, in fact they use this form as their standards

OpenStudy (anonymous):

Duh! I get it. thanks!

OpenStudy (amistre64):

your prolly used to "simplifying" everything; math works best when you just use it to make life simplier instead ;)

OpenStudy (anonymous):

figuring out how to simplify... can sometimes be complicated

OpenStudy (anonymous):

i understand.. it threw me back to my precal days i get it.. although i need help graphing.. its been a while.. 20/3 and 20/5 how do i graph

OpenStudy (anonymous):

You need to define points. y = 4 when x = 0 y = 0 when x = 20/3

OpenStudy (anonymous):

ohh i see okay thanks